2x+y=-7and y= 3x+3 I HAVE TO FIND OUT ANSWER FOR X AND Y
Answer by mananth(16946)

2 x + 1 y = -7 .............1
-3 x + 1 y = 3 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-2 x -1 y = 7
-3 x 1 y = 3
Add the two equations
-5 x = 10
/ -5
x = -2
plug value of x in (1)
2 x + 1 y = -7
-4 + 1 y = -7
1 y = -7 + 4
1 y = -3
y = -3
(-2,-3)